OpenStudy (anonymous):

can you please solve 5 and 10

OpenStudy (anonymous):

@dan815

OpenStudy (muggy21):

What?

OpenStudy (ibrafacts):

cant see problem try again

OpenStudy (muggy21):

I'm not seeing a question here

OpenStudy (anonymous):

Find the geometric mean between each pair of numbers

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@pooja195

OpenStudy (muggy21):

7.07107?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (freckles):

what is the question? I see you have asked to find the geometric mean between each pair of numbers... but you have not given any pair of numbers yet... the geometric mean of a and b is given below though: \[\sqrt{a \cdot b}\]

OpenStudy (anonymous):

Okay so how do i solve it

OpenStudy (freckles):

I'm saying I cannot help you solve anything because you have no pair of numbers given...

OpenStudy (anonymous):

5 and 10

OpenStudy (freckles):

use the definition I gave you above to find the geometric mean of two numbers \[\text{ geometric mean of } a \text{ and } b \text{ is } \sqrt{ a \cdot b}\] just replace a with 5 and b with 10

OpenStudy (anonymous):

\[\sqrt{50}\]

OpenStudy (freckles):

right notice there is a perfect square factor in 50

OpenStudy (freckles):

50=25*2 so you can rewrite the sqrt(50) as...

OpenStudy (anonymous):

25

OpenStudy (freckles):

yes 25 is a perfect square

OpenStudy (freckles):

sqrt(25)=?

OpenStudy (anonymous):

5

OpenStudy (freckles):

\[\sqrt{50}=\sqrt{25} \cdot \sqrt{2}=5 \cdot \sqrt{2}=5 \sqrt{2}\]

OpenStudy (anonymous):

wow that is easy i think i know it now